# Various Properties of Various Ultrafilters, Various Graph Width Parameters, and Various Connectivity Systems

### v6

@@ Abstract
Filters are collections of sets that are closed under supersets and finite intersections, serving as fundamental tools in topology and set theory. An ultrafilter, a maximal filter on a set, plays a crucial role in these fields by rigorously handling limits, convergence, and compactness. A connectivity system is defined as a pair $(X,f)$, where $X$ is a finite set and $f$ is a symmetric submodular function. Understanding the duality in these parameters elucidates the relationship between different decompositions and measures of a graph’s complexity. In this paper, we delve into ultrafilters on connectivity systems, applying Tukey’s Lemma to these systems. Additionally, we explore prefilters, ultra-prefilters, and subbases within the context of connectivity systems. Furthermore, we introduce and investigate new parameters related to width, length, and depth, enhancing our understanding of these mathematical structures. We also investigate and compare various graph width parameters and their related parameters. (It also includes information similar to that found in surveys, aiming to promote future research on graph width parameters.)
@@ 1.1 Filter and Graph width parameters
Filters are essential collections of sets in topology and set theory, characterized by their closure under supersets and finite intersections. They can be viewed as a way to "focus" on certain subsets of a space, much like a lens that sharpens specific details while filtering out others. An ultrafilter, being a maximal filter on a set, is particularly significant for addressing fundamental concepts such as limits, convergence, and compactness. Its unique properties render it indispensable in various fields, including non-standard analysis, model theory, social choice theory (social judgments), group theory, boolean algebra, geometry, probability theory, vector theory, semigroup theory, abstract algebra, topology, set theory, infinite combinatorics, fuzzy theory, graph theory, matroid theory, lattice theory, and first-order logic, where it provides powerful tools for both mathematical and logical applications[304, 302, 719, 389, 3, 167, 340, 475, 642, 78, 144, 544, 295, 142, 297, 200].
Graph theory, a key branch of mathematics, delves into the study of networks consisting of nodes and edges, with a focus on their paths, structures, and properties [696, 187, 99]. Among the critical metrics in this field is the "graph width parameter," which measures the maximum width across all cuts or layers within a hierarchical decomposition of the graph. This metric is crucial for analyzing a graph’s complexity and structure, serving as a primary factor in transforming computationally hard graph problems into more tractable ones when the graph class is restricted to having bounded width. Various width parameters have been rigorously explored and are widely recognized in the literature [330, 577, 232, 242, 304, 19, 613, 486, 370, 423, 279, 485, 405, 296, 667, 298]. For further information, see Appendix A, “Various Width Parameters,” and Appendix C, “Comparing Various Graph Parameters (Over 70 Parameters).”
@@ 1.2 Our Contribution
  • Section 6: We investigate new parameters such as width, length, and depth, as well as the Ultrafilter game on connectivity systems, ultraproduct on connectivity systems, and the Axiom of Choice on connectivity systems. We also consider a Small Set Expansion [702, 39, 714, 137, 251, 70].
@@
Note 1.1.
@@ 2.1.2 Notation for undirected graph
A path graph is a type of tree that consists of a sequence of vertices where each vertex is connected to exactly two others, except for the two endpoints, which are connected to only one other vertex. A linear ordering of a graph’s vertices is an arrangement of the vertices in a linear sequence in such a way that each vertex comes before all vertices to which it is connected by edges in the sequence, except possibly the previous vertex.
@@
A tree is a connected, acyclic graph where any two vertices are connected by exactly one path, symbolizing hierarchical relationships. The root of a tree is the topmost node from which all other nodes descend. In a tree, vertices with a degree of 1 are called leaves, while all other vertices are referred to as inner vertices. A ternary tree is a specific type of tree where all inner vertices have a degree of 3.
@@ 2.2 Symmetric Submodular Function and Connectivity System
2.3 Ultrafilter on Connectivity System
In this subsection. we consider about ultrafilter on Connectivity System. First, we consider about ultrafilter for set theory.
2.3.1 Ultrafilter for set theory
First, let’s introduce the general concept of an Ultrafilter in set theory. The definition of an Ultrafilter in Set Theory is described as follows.
@@ 2.3.1 Ultrafilter for set theory
Let $X$ be a set. A collection $\mathcal{F}\subseteq 2^{X}$ is called a filter on $X$ if it satisfies the following conditions111*1:
  1. (B1)
@@
Another example of a filter is the Fréchet filter on an infinite set $X$. The Fréchet filter consists of all cofinite subsets of $X$ (i.e., subsets whose complements are finite). This collection forms a filter because:
@@
Several related concepts to Filters and Ultrafilters have been proposed. Here, I will introduce just one: the Weak Filter. This concept, discussed in the field of logic, is used to interpret defaults through a generalized "most" quantifier in first-order logic [476, 621, 34, 477, 141]. The definition and an example are provided below.
@@
2.3.2 Ultrafilter on Connectivity system
We introduce some properties of the ultrafilter on the connectivity system $(X,f)$ as an extension of the Ultrafilter on Boolean Algebras. First, we introduce definitions of a filter and an ultrafilter on a connectivity system [304]. These concepts extend the traditional notions of filter and ultrafilter from set theory by incorporating the condition of symmetric submodularity. Also, an ultrafilter on the connectivity system $(X,f)$ is co-Maximal ideal [713] on the connectivity system $(X,f)$.

### v8

@@ Abstract
This paper investigates ultrafilters in the context of connectivity systems, defined as pairs $(X,f)$ where $X$ is a finite set and $f$ is a symmetric submodular function. Ultrafilters, essential in topology and set theory, are extended to these systems, with a focus on their relationship to graph width parameters, which help analyze graph complexity.
We demonstrate theorems for ultrafilters on connectivity systems and explore related concepts such as prefilters, ultra-prefilters, and subbases. New parameters for width, length, and depth are introduced, providing further insight into graph width. The study also includes a comparison of various graph width parameters and their related concepts, offering a foundation for future research in graph theory and computational complexity. Additionally, we explore connections to other mathematical disciplines, including set theory, lattice theory, and matroid theory, expanding the scope of ultrafilters and graph width. (It also includes information similar to that found in surveys, aiming to promote future research on graph width parameters.)
@@ 1.1 Filter and Graph width parameters
Filters are essential collections of sets in topology and set theory, characterized by their closure under supersets and finite intersections. They can be viewed as a way to "focus" on certain subsets of a space, much like a lens that sharpens specific details while filtering out others. An ultrafilter, being a maximal filter on a set, is particularly significant for addressing fundamental concepts such as limits, convergence, and compactness. Its unique properties render it indispensable in various fields, including non-standard analysis, model theory, social choice theory (social judgments), group theory, boolean algebra, geometry, probability theory, vector theory, semigroup theory, abstract algebra, topology, set theory, infinite combinatorics, fuzzy theory, graph theory, matroid theory, lattice theory, and first-order logic, where it provides powerful tools for both mathematical and logical applications[529, 526, 1240, 657, 490, 493, 6, 298, 579, 795, 1109, 156, 255, 912, 514, 253, 517, 354].
Graph theory, a key branch of mathematics, delves into the study of networks consisting of nodes and edges, with a focus on their paths, structures, and properties [1207, 338, 183]. Among the critical metrics in this field is the "graph width parameter," which measures the maximum width across all cuts or layers within a hierarchical decomposition of the graph. This metric is crucial for analyzing a graph’s complexity and structure, serving as a primary factor in transforming computationally hard graph problems into more tractable ones when the graph class is restricted to having bounded width. Various width parameters have been rigorously explored and are widely recognized in the literature [565, 966, 410, 427, 529, 48, 1038, 814, 618, 711, 500, 813, 496, 499, 510, 683, 515, 1153, 518]. For further information, see Appendix A, “Various Width Parameters,” and Appendix C, “Comparing Various Graph Parameters (Over 70 Parameters).”
@@ 1.2 Our Contribution
  • Section 6: We investigate new many parameters such as width, length, and depth, as well as the Ultrafilter game on connectivity systems, ultraproduct on connectivity systems, and the Axiom of Choice on connectivity systems. We also consider a Small Set Expansion [1217, 93, 1232, 241, 443, 142].
@@
We aim to make new contributions to the study of Graph Width Parameters and Graph Algorithms by interpreting key concepts like Tree-decomposition, Branch-decomposition, Tangles, and their related graph parameters through the lens of various mathematical disciplines. For instance, in this paper, we explore concepts from Set Theory (Ultrafilter, prefilter), Lattice Theory (chain, antichain), Model Theory (Ultraproduct), Hypergraph Theory (fuzzy hypergraph, clutter), Matroid Theory (Single-element-extension), Topology (P-point, Q-point), Fuzzy Theory (Fuzzy Tree-width, bipolar fuzzy graphs), First-order Logic (weak ultrafilter), Social Judgment (Voting system, Quasi Ultrafilter), Game Theory (cops and robbers, simple game, monotone search game), and Sperner Theory (Sperner system and trace), among others.
@@ 2.1.1 Notation for Set theory
Example 2.3 (Powerset of a Set).
Let $A=\{1,2\}$. The powerset of $A$, denoted as $2^{A}$, is the set of all possible subsets of $A$. This includes the empty set, the individual elements, and the set $A$ itself. Therefore, the powerset of $A$ is:
$2^{A}=\{\emptyset,\{1\},\{2\},\{1,2\}\}.$
Similarly, for a set $B=\{a,b,c\}$, the powerset $2^{B}$ is:
$2^{B}=\{\emptyset,\{a\},\{b\},\{c\},\{a,b\},\{a,c\},\{b,c\},\{a,b,c\}\}.$
In this case, there are $82^{3}=$ 8 subsets, as the size of the powerset is $2^{|B|}$, where $|B|$ is the number of elements in the original set.
@@
Example 2.5 (Partition of a Set).
Consider the set $X=\{1,2,3,4,5,6\}$. A partition of $X$ is a collection of non-overlapping, non-empty subsets whose union equals $X$. For instance, the following subsets form a partition of $X$:
$P=\{\{1,2\},\{3,4\},\{5,6\}\}.$
Here, each element of $X$ belongs to exactly one subset, and the subsets do not overlap. Hence, $P$ is a valid partition of $X$.
Definition 2.6.
In set theory, a separation of a set $X$ is a partition of $X$ into two disjoint, non-empty subsets $A$ and $B$ such that their union equals $X$ (i.e., $A\cup B=X$ and $A\cap B=\emptyset$). The subsets $A$ and $B$ are often referred to as the separated components of $X$.
Definition 2.7.
In set theory, the cardinality of a set refers to the number of elements in the set. For finite sets, the cardinality is simply the count of elements in the set.
Formally, the cardinality of a set $A$ is denoted by $|A|$. If there exists a bijection (a one-to-one correspondence) between two sets $A$ and $B$, they are said to have the same cardinality, i.e., $|A|=|B|$.
@@ 2.1.2 Notation for undirected graph
(cf.[218]) A path graph is a type of tree that consists of a sequence of vertices where each vertex is connected to exactly two others, except for the two endpoints, which are connected to only one other vertex. A linear ordering of a graph’s vertices is an arrangement of the vertices in a linear sequence in such a way that each vertex comes before all vertices to which it is connected by edges in the sequence, except possibly the previous vertex.
@@
Definition 2.11 (Connectedness).
(cf.[1175]) A graph $G=(V,E)$ is said to be connected if for every pair of vertices $u,v\in V$, there exists a path $P\subseteq G$ that connects $u$ and $v$. Formally, $G$ is connected if:
$\forall u,v\in V,\exists P\subseteq G\text{ such that }P\text{ is a path from }u\text{ to }v.$
@@
(cf.[203]) A tree is a connected, acyclic graph where any two vertices are connected by exactly one path, symbolizing hierarchical relationships. The root of a tree is the topmost node from which all other nodes descend. In a tree, vertices with a degree of 1 are called leaves, while all other vertices are referred to as inner vertices. A ternary tree is a specific type of tree where all inner vertices have a degree of 3.
@@
Definition 2.18.
A clique in an undirected graph $G=(V,E)$ is a subset $C\subseteq V$ of vertices such that every pair of distinct vertices in $C$ is connected by an edge in $G$. Formally, for all $u,v\in C$ with $u\neq v$, the edge $(u,v)\in E$.
Example 2.19.
Consider a graph $G$ with vertex set $V(G)=\{v_{1},v_{2},v_{3},v_{4},v_{5}\}$ and edge set
$E(G)=\{\{v_{1},v_{2}\},\{v_{1},v_{3}\},\{v_{1},v_{4}\},\{v_{2},v_{3}\},\{v_{2},v_{4}\},\{v_{3},v_{4}\},\{v_{4},v_{5}\}\}.$
The subset $C=\{v_{1},v_{2},v_{3},v_{4}\}$ forms a clique because every pair of distinct vertices in $C$ is connected by an edge:
  • $\{v_{1},v_{2}\},\{v_{1},v_{3}\},\{v_{1},v_{4}\}\in E(G)$,
  • $\{v_{2},v_{3}\},\{v_{2},v_{4}\}\in E(G)$,
  • $\{v_{3},v_{4}\}\in E(G)$.
However, the subset $\{v_{3},v_{4},v_{5}\}$ is not a clique, as $\{v_{3},v_{5}\}\notin E(G)$.
Definition 2.20 (union).
(cf.[357, 503]) The union of two graphs $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ is a graph $G=(V,E)$ where:
  • The vertex set $V$ is the union of the vertex sets of $G_{1}$ and $G_{2}$:
    $V=V_{1}\cup V_{2}.$
  • The edge set $E$ is the union of the edge sets of $G_{1}$ and $G_{2}$:
    $E=E_{1}\cup E_{2}.$
Thus, the union of $G_{1}$ and $G_{2}$ combines the vertices and edges of both graphs, without duplicating any elements.
Definition 2.21 (Intersection).
(cf.[885, 289]) The intersection of two graphs $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ is a graph $G=(V,E)$ where:
  • The vertex set $V$ is the intersection of the vertex sets of $G_{1}$ and $G_{2}$:
    $V=V_{1}\cap V_{2}.$
  • The edge set $E$ is the intersection of the edge sets of $G_{1}$ and $G_{2}$:
    $E=E_{1}\cap E_{2}.$
Thus, the intersection of $G_{1}$ and $G_{2}$ consists of only the vertices and edges that are common to both graphs.
@@ 2.1.3 Notation for directed graph
Example 2.25 (Directed Tree with Specific Values).
Consider the directed tree $T=(V,E)$, where the set of vertices is $V=\{1,2,3,4,5\}$ and the set of directed edges is $E=\{(1,2),(1,3),(2,4),(2,5)\}$. The directed tree can be illustrated as follows: $1\rightarrow 2\rightarrow 4$, $1\rightarrow 2\rightarrow 5$, $1\rightarrow 3$.
In this example:
  • The root of the tree is vertex $1$, as it has no incoming edges.
  • The vertices $4$ and $5$ are leaves, as they have exactly one incoming edge and no outgoing edges.
  • The tree has $5$ vertices and $4$ directed edges, satisfying the condition that a directed tree with $n$ vertices has $n-1$ edges.
  • The tree is connected, acyclic, and contains no directed cycles, making it a valid directed tree.
This is also an example of a subcubic directed tree, as no vertex has a degree greater than 3. For instance, vertex $2$ has an in-degree of 1 and an out-degree of 2, giving it a total degree of 3.
@@ 2.2 Symmetric Submodular Function and Connectivity System
2.3 Ultrafilter for set theory
We consider about ultrafilter on Connectivity System. First, let’s introduce the general concept of an Ultrafilter in set theory. The definition of an Ultrafilter in Set Theory is described as follows.
@@ 2.3 Ultrafilter for set theory
Let $X$ be a set. A collection $\mathcal{F}\subseteq 2^{X}$ is called a filter on $X$ if it satisfies the following conditions:
  1. (BF1)
@@
Definition 2.35.
(cf.[745, 50]) Given a filter $\mathcal{F}$ on a set $X$, we say a subset $A\subseteq X$ is $\mathcal{F}$-stationary if, for all $B\in\mathcal{F}$, we have $A\cap B\neq\emptyset$.
@@
(cf.[562, 713]) Example of a filter is the Fréchet filter on an infinite set $X$. The Fréchet filter consists of all cofinite subsets of $X$ (i.e., subsets whose complements are finite). This collection forms a filter because:
@@
Example 2.38 (neighbourhood filter).
(cf.[435, 692]) Let $X$ be a topological space and let $x\in X$ be a point. The collection $\mathcal{N}(x)$, defined as the set of all neighbourhoods of $x$, forms a filter on $X$. This is because:
  1. 1.
    For any neighbourhoods $N_{1},N_{2}\in\mathcal{N}(x)$, their intersection $N_{1}\cap N_{2}$ is also a neighbourhood of $x$, thus $N_{1}\cap N_{2}\in\mathcal{N}(x)$
  2. 2.
    If $N\in\mathcal{N}(x)$ and $N\subseteq N^{\prime}$, where $N^{\prime}$ is an open set containing $x$, then $N^{\prime}\in\mathcal{N}(x)$
  3. 3.
    $x$ belongs to every neighbourhood in $\mathcal{N}(x)$, ensuring that $\mathcal{N}(x)$ is non-empty.
Example 2.39.
Consider the set $X=\mathbb{R}$, the set of all real numbers. The collection =$\mathcal{F}=\{A\subseteq\mathbb{R}\mid\exists M\in\mathbb{R},\ A\supseteq[M,\infty)\}$ is a filter on $\mathbb{R}$. This collection forms a filter because:
  • The intersection of any two sets that contain some $[M,\infty)$ also contains a set of the form $[M^{\prime},\infty)$ for some $M^{\prime}\geq\max\{M_{1},M_{2}\}$, so it is in $\mathcal{F}$.
  • If $A\in\mathcal{F}$ and $A\subseteq B\subseteq\mathbb{R}$, then $B\in\mathcal{F}$.
  • The empty set does not contain any interval of the form $[M,\infty)$, so it is not in the filter.
Example 2.40.
Consider a topological space $(X,\tau)$. The collection of all dense open sets  [^1] Dense open sets are subsets of a topological space that are both open and dense, meaning they intersect every non-empty open set, covering the entire space in a ”thick” way(cf.[916, 593]). in $X$ forms a filter. This is because the intersection of two dense open sets is also a dense open set, and any superset of a dense open set is dense and open. The empty set is not dense, so it is not in the filter.
Example 2.41.
In a measure space $(X,\mathcal{M},\mu)$, where $\mu$ is a measure on the sigma-algebra $\mathcal{M}$  [^2] A sigma-algebra is a collection of subsets of a given set that is closed under complements, countable unions, and countable intersections. It provides the framework for defining measures in measure theory., the collection of all subsets of $X$ that have positive measure (i.e., $\mu(A)>0$) forms a filter. This filter includes all sets with non-zero measure, and satisfies the filter conditions.
Example 2.42 (Voting system).
 [^3] A voting system is a method by which voters select candidates or policies, usually involving preferences or choices aggregated to determine the final outcome, often using mechanisms like majority, plurality, or ranked voting. (cf.[602, 95, 292, 861])
Consider a set $X$ that represents a collection of voters $X=\{v_{1},v_{2},\dots,v_{n}\}$. Each voter has a preference for a particular candidate in an election. We define a collection $\mathcal{F}$ of subsets of $X$ as follows :
  • A subset $A\subseteq X$ belongs to $\mathcal{F}$ if the majority of voters in $A$ support a particular candidate, say Candidate $C$.
This collection $\mathcal{F}$ satisfies the properties of an ultrafilter:
  1. 1.
    Non-emptiness and Completeness: The set $\mathcal{F}$ is non-empty because there exists at least one subset of voters (e.g., the entire set $X$) that supports Candidate $C$.
  2. 2.
    Closure under Intersection: If two subsets $A$ and $B$ are in $\mathcal{F}$, meaning that the majority in both $A$ and $B$ supports Candidate $C$, then their intersection $A\cap B$ also belongs to $\mathcal{F}$ because the majority in $A\cap B$ still supports Candidate $C$.
  3. 3.
    Closure under Supersets: If $A$ is in $\mathcal{F}$ and $A\subseteq B\subseteq X$, then $B$ is also in $\mathcal{F}$. This is because if the majority of $A$ supports Candidate $C$, and $B$ includes all of $A$, then the majority in $B$ still supports $C$.
  4. 4.
    Decisiveness: For any subset $A\subseteq X$, either $A$ or its complement $X\setminus A$ is in $\mathcal{F}$. This corresponds to the idea that, in any group of voters, either the majority supports $C$ or the majority does not, with no undecided or neutral outcomes.
This voting system example illustrates an ultrafilter by capturing the idea of a decisive majority in a set of voters. In this context, the ultrafilter $\mathcal{F}$ helps identify the subsets of voters that consistently support a particular candidate across various possible groupings.
Several related concepts to Filters and Ultrafilters have been proposed. Here, we introduce just one: the Weak Filter. This concept, discussed in the field of logic, is used to interpret defaults through a generalized "most" quantifier in first-order logic [796, 798, 1066, 1064, 797, 1063, 1062, 86, 799, 252]. The definition and an example are provided below.
@@
Several properties of ultrafilters in set theory are listed below. As will be explained later, in a connectivity system, the focus is on sets where $f(A)\leq k$. Therefore, the properties of ultrafilters in set theory do not necessarily hold in a connectivity system.
Definition 2.46 (Finite intersection property).
(cf.[673]) Let $\mathcal{C}$ be a collection of subsets of $X$. We say that $\mathcal{C}$ has the finite intersection property (FIP) if, for any $D_{1},\dots,D_{n}\in\mathcal{C}$, we have $D_{1}\cap\cdots\cap D_{n}\neq\emptyset$.
Theorem 2.47.
Let $X$ be a finite set, and let $\mathcal{F}$ be a filter on $X$ with $A\subseteq X$. Then, $\mathcal{F}\cup\{A\}$ has the finite intersection property (FIP) if and only if $X\setminus A\notin\mathcal{F}$.
Proof 2.48.
Let $X$ be a finite set, $\mathcal{F}$ be a filter on $X$, and $A\subseteq X$. We aim to prove that $\mathcal{F}\cup\{A\}$ has the finite intersection property (FIP) if and only if $X\setminus A\notin\mathcal{F}$.
Assume that $\mathcal{F}\cup\{A\}$ has the finite intersection property. By definition, this means that for any finite collection of sets $D_{1},D_{2},\dots,D_{n}\in\mathcal{F}\cup\{A\}$, we have $D_{1}\cap D_{2}\cap\dots\cap D_{n}\neq\emptyset$. Now, suppose for contradiction that $X\setminus A\in\mathcal{F}$. Consider the two sets $A$ and $X\setminus A$. Clearly, $A\cap(X\setminus A)=\emptyset$. Since $A\in\mathcal{F}\cup\{A\}$ and $X\setminus A\in\mathcal{F}$, their intersection should be non-empty if $\mathcal{F}\cup\{A\}$ has the FIP. But $A\cap(X\setminus A)=\emptyset$, which contradicts the assumption that $\mathcal{F}\cup\{A\}$ has the FIP. Therefore, it must be that $X\setminus A\notin\mathcal{F}$.
Now assume that $X\setminus A\notin\mathcal{F}$. We need to show that $\mathcal{F}\cup\{A\}$ has the finite intersection property. Consider any finite collection of sets $D_{1},D_{2},\dots,D_{n}\in\mathcal{F}\cup\{A\}$. If none of these sets are equal to $A$, then they all belong to $\mathcal{F}$. Since $\mathcal{F}$ is a filter, we know that $D_{1}\cap D_{2}\cap\dots\cap D_{n}\neq\emptyset$ because filters have the FIP by definition. If one of the sets is $A$, then the intersection becomes $A\cap D_{1}\cap D_{2}\cap\dots\cap D_{n}$, where all other sets are in $\mathcal{F}$. Since $X\setminus A\notin\mathcal{F}$, $A\cap D_{i}\neq\emptyset$ for all $D_{i}\in\mathcal{F}$, meaning that $A\cap D_{1}\cap D_{2}\cap\dots\cap D_{n}\neq\emptyset$. Thus, $\mathcal{F}\cup\{A\}$ has the FIP.
We have shown that $\mathcal{F}\cup\{A\}$ has the finite intersection property if and only if $X\setminus A\notin\mathcal{F}$.
Definition 2.49 (uniform ultrafilter).
(cf.[937]) An ultrafilter $\mathcal{U}$ on $Y$ is called uniform if $|A|=|Y|$ for every $A\in\mathcal{U}$.
Theorem 2.50.
Let $\mathcal{U}$ be an ultrafilter on $Y$, and suppose that $Z\in\mathcal{U}$ has minimal cardinality among the sets in $\mathcal{U}$. Then $\mathcal{U}\cap Z$ is a uniform ultrafilter on $Z$.
Proof 2.51.
We need to show that $\mathcal{U}\cap Z$ is a uniform ultrafilter on $Z$.
Since $\mathcal{U}$ is an ultrafilter on $Y$ and $Z\in\mathcal{U}$, the family $\mathcal{U}\cap Z$ consists of subsets of $Z$ that are in $\mathcal{U}$. To verify that $\mathcal{U}\cap Z$ is an ultrafilter on $Z$, we must confirm that it satisfies the properties of an ultrafilter: - For any set $A\subseteq Z$, either $A\in\mathcal{U}\cap Z$ or $Z\setminus A\in\mathcal{U}\cap Z$. This holds because $Z\in\mathcal{U}$, and $\mathcal{U}$ is an ultrafilter on $Y$, meaning for any subset of $Z$, either it or its complement (within $Z$) is in $\mathcal{U}$. - $\mathcal{U}\cap Z$ contains $Z$ itself, which ensures non-emptiness.
Thus, $\mathcal{U}\cap Z$ is an ultrafilter on $Z$.
By assumption, $Z$ has minimal cardinality among the sets in $\mathcal{U}$. For $\mathcal{U}\cap Z$ to be a uniform ultrafilter on $Z$, we need to show that $|A|=|Z|$ for every $A\in\mathcal{U}\cap Z$.
Since $Z$ is of minimal cardinality, any subset $A\in\mathcal{U}\cap Z$ must be such that $|A|=|Z|$. If $|A|<|Z|$ for some $A\in\mathcal{U}\cap Z$, then $A$ would have a smaller cardinality than $Z$, contradicting the assumption that $Z$ has minimal cardinality in $\mathcal{U}$.
Thus, every set in $\mathcal{U}\cap Z$ has the same cardinality as $Z$, meaning that $\mathcal{U}\cap Z$ is a uniform ultrafilter on $Z$.
We have shown that $\mathcal{U}\cap Z$ is both an ultrafilter and uniform on $Z$, completing the proof.
Theorem 2.52 (ultrafilter on $A$ induced by $\mathcal{U}$).
Let $\mathcal{U}$ be an ultrafilter on $X$ and let $A\in\mathcal{U}$. Define the set
$A\cap\mathcal{U}:=\{A\cap B:B\in\mathcal{U}\}.$
Then $A\cap\mathcal{U}$ is an ultrafilter on $A$.
Proof 2.53.
Since $\mathcal{U}$ is an ultrafilter on $X$, it is non-empty and closed under intersections. For any $B_{1},B_{2}\in\mathcal{U}$, we have $A\cap B_{1}\in A\cap\mathcal{U}$ and $A\cap B_{2}\in A\cap\mathcal{U}$. Therefore,
$(A\cap B_{1})\cap(A\cap B_{2})=A\cap(B_{1}\cap B_{2})\in A\cap\mathcal{U}.$
Thus, $A\cap\mathcal{U}$ is non-empty and closed under intersections.
If $A\cap B\in A\cap\mathcal{U}$ and $A\cap B\subseteq A\cap C$ for some $C\subseteq X$, then $B\subseteq C$ since $\mathcal{U}$ is an ultrafilter on $X$. Thus, $A\cap C\in A\cap\mathcal{U}$, satisfying the superset condition.
For any subset $C\subseteq A$, we need to show that either $C\in A\cap\mathcal{U}$ or $A\setminus C\in A\cap\mathcal{U}$. Since $\mathcal{U}$ is an ultrafilter on $X$, for any subset $C\subseteq A$, either $C\in\mathcal{U}$ or $X\setminus C\in\mathcal{U}$. - If $C\in\mathcal{U}$, then $A\cap C\in A\cap\mathcal{U}$. - If $X\setminus C\in\mathcal{U}$, then $A\cap(X\setminus C)=A\setminus C\in A\cap\mathcal{U}$.
Thus, either $C\in A\cap\mathcal{U}$ or $A\setminus C\in A\cap\mathcal{U}$, satisfying the maximality condition.
We have shown that $A\cap\mathcal{U}$ is non-empty, closed under intersections, contains supersets, and satisfies the maximality condition. Therefore, $A\cap\mathcal{U}$ is an ultrafilter on $A$, called the ultrafilter on $A$ induced by $\mathcal{U}$.
@@
2.4 Ultrafilter on Connectivity system
We introduce some properties of the ultrafilter on the connectivity system $(X,f)$ as an extension of the Ultrafilter on Boolean Algebras.
First, we consider ultrafilters on graphs. This concept extends the idea of ultrafilters from set theory to graphs and has a dual relationship with tree-width[495]. Refer to Appendix A for the definition of tree-width.
Definition 2.54.
[495] Let $G$ be a graph. A G-Ultrafilter of order $k$ is a family $\mathcal{F}$ of separations of $G$ satisfying the following conditions:
  • (FG0)
    The order of all separations $(A,B)\in\mathcal{F}$ is less than $k$.
  • (FG1)
    For all separations $(A,B)$ of $G$ with order less than $k$, either $(A,B)\in\mathcal{F}$ or $(B,A)\in\mathcal{F}$.
  • (FG2)
    If $(A_{1},B_{1})\in\mathcal{F}$, $A_{1}\subseteq A_{2}$, and $(A_{2},B_{2})$ is a separation of $G$ with order less than $k$, then $(A_{2},B_{2})\in\mathcal{F}$.
  • (FG3)
    If $(A_{1},B_{1})\in\mathcal{F}$ and $(A_{2},B_{2})\in\mathcal{F}$, and $(A_{1}\cap A_{2},B_{1}\cup B_{2})$ is a separation of $G$ with order less than $k$, then $(A_{1}\cap A_{2},B_{1}\cup B_{2})\in\mathcal{F}$.
  • (FG4)
    If $V(A)=V(G)$, then $(A,B)\in\mathcal{F}$.
Next, we introduce definitions of a filter and an ultrafilter on a connectivity system [529]. These concepts extend the traditional notions of filter and ultrafilter from set theory by incorporating the condition of symmetric submodularity. Also, an ultrafilter on the connectivity system $(X,f)$ is co-Maximal ideal [1231] on the connectivity system $(X,f)$.